# MMGT: Motion Mask Guided Two-Stage Network for Co-Speech Gesture Video Generation

### v1

@@ III-B Motion feature generation
where $p^{(0)}$ represents the static pose keypoints of the subject in the reference image $I_{0}$. To introduce variations for motion modeling, Gaussian noise is added to the ground-truth pose sequence $P_{\text{GT}}$ to obtain noisy motion representations,
$x^{\prime}_{t}=P_{\text{GT}}+\mathcal{N}(0,\sigma^{2}),$ (4)
where $x^{\prime}_{t}$ represents the noisy pose at time step $t$. The initial pose $p^{(0)}$ is repeated along the temporal dimension and added to $x_{t}$,
$x_{t}=x^{\prime}_{t}+\mathcal{R}(p^{(0)}).$ (5)
where $\mathcal{R}(\cdot)$ denotes the delayed time dimension replication N copies. Motion features have been categorized into facial expression motion features and gesture motion features as
@@
where $C_{f}$ corresponds to the keypoints relevant to head motion, while $C_{b}$ corresponds to the keypoints for gesture motion. To apply masking to $x_{t}$ along the feature dimension $C$, we construct the mask $M^{f}$ as
@@
Next, the complementary mask $M^{b}$ is obtained as $M^{b}=1-M^{f}$. These masks are then applied to the input $x_{t}$ to isolate the respective feature representations, as defined below
@@
where $x_{t}^{f}$ and $x_{t}^{b}$ correspond to the features for head and gesture pose sequence, respectively. Distinct encoders are employed to process these input conditions, transforming them into separate feature embeddings
@@
where $f_{a}$ represents the audio feature embedding, with $E_{a}$ being a pre-trained audio encoder [11], while $E_{f}$ and $E_{b}$ are encoders for the face and body motion features, $f_{f}$ and $f_{b}$, respectively. Finally, the output of the first stage is expressed as
@@
where $t$ represents the temporal embedding. These features are subsequently fed into the Spatial Mask-Guided Audio Pose Generation (SMGA) Network , denoted as $D_{\text{SMGA}}$, which decodes them to produce a sequence of audio-driven poses $\hat{x}_{0}$. The generated pose sequence $\hat{x}_{0}$ is then visualized as the pose video $V_{P}$.
Motion Mask Generation. As depicted in Algorithm 1, our approach introduces a novel method for generating motion masks based on keypoints, represented as $P\in\mathbb{R}^{T\times C_{p}}$. Here, $C_{p}=C_{f}+C_{h}+C_{l}$ corresponds to face, hand gestures, and lips, respectively. For each frame within a video sequence, our algorithm transforms normalized keypoint coordinates into pixel coordinates and calculates bounding boxes for designated sections. These bounding boxes are then utilized to construct binary masks, setting active regions to a value of 255. The generated masks include $V^{f}_{M}$ for face, $V_{M}^{l}$ for limb movements, and $V_{M}^{h}$ for hand gestures. Notably, as the head and hand masks are non-overlapping, they are combined into a single mask, denoted as $V_{M}^{f+h}$, and we set the background mask to $V_{M}^{b}=(255-V_{M}^{f+h})$.
Details of the SMGA Network Structure. The SMGA network is carefully designed to capture the temporal and semantic relationships between motion features and audio input. As shown in Fig. 4, its architecture consists of two motion blocks and a merging module. The two motion blocks process the input features $f_{f}$ or $f_{b}$ that represent different body parts separately through a combination of self-attention, FiLM (Feature-wise Linearly Module), and cross-attention mechanisms to capture the correlation between different body parts and audio. Among them, the self-attention mechanism is used to capture the correlation of the motion features themselves, and the calculation formula is as
where $Q_{f}$, $K_{f}$, $V_{f}$ are derived from motion features ($f_{f}$ or $f_{b}$). Then, the cross-attention mechanism in different motion modules will capture the different correlations between different body parts and audio embeddings. The specific formulas are as
where $Q_{f}$ is derived from motion features ($f_{f}$ or $f_{b}$), and $K_{f_{a}}$, $V_{f_{a}}$ are derived from the audio features $f_{a}$. Through this distraction mechanism, we achieve coherent and coordinated speech actions for different semantic regions (e.g., body and face). In addition, the introduced FiLM layer plays a key role by dynamically adjusting the action features according to the audio embedding, thereby facilitating precise cross-modal alignment.
To ensure the coordination of head movements and hand gestures, the network employs a fusion strategy. First, a multi-layer perceptron (MLP) combines the features of the head movements and hand gestures. Then, a FiLM layer is used to dynamically adapt the fused features. This process is mathematically expressed as
$f_{\text{merged}}=\text{FiLM}(\text{MLP}(f_{f}+f_{b})),$ (13)
where $f_{f}$ and $f_{b}$ represent the head and gesture motion features, respectively. The final dynamic feature map, $f_{\text{merged}}$, is then used to predict the pose sequence $\hat{x}_{0}$, which corresponds to the audio-driven pose sequence. Through this multi-step integration, the model can generate a pose sequence that is highly consistent with the input audio and motion features, and is also temporally coherent and semantically rich.
Figure 5: The details of inputs during model training.
@@
0:  Pre-keypoints $\hat{P}\in\mathbb{R}^{N\times{C_{p}}},{C_{p}}=(C_{f}+C_{h}+C_{l})$
@@
Loss Function. The training process optimizes the SMGA network using a multi-component loss function to ensure accurate motion generation. The reconstruction loss is defined as
@@
where $x_{t}$ and $\hat{x}_{t}$ represent the ground-truth and predicted motion features, respectively. The velocity loss is
$\mathcal{L}_{\text{vel}}=\frac{1}{T-1}\sum_{t=1}^{T-1}\|(x_{t+1}-x_{t})-(\hat{x}_{t+1}-\hat{x}_{t})\|_{2}^{2}.$ (15)
@@
$\quad-\left(\hat{x}_{t+2}-2\hat{x}_{t+1}+\hat{x}_{t}\right)\Big{\|}_{2}^{2}.$ (16)
Each loss term for head and gesture motion can be defined as
$\mathcal{L}_{f}=\mathcal{L}_{\text{rec}}^{f}+\mathcal{L}_{\text{vel}}^{f}+\mathcal{L}_{\text{acc}}^{f},\quad\mathcal{L}_{b}=\mathcal{L}_{\text{rec}}^{b}+\mathcal{L}_{\text{vel}}^{b}+\mathcal{L}_{\text{acc}}^{b}.$ (17)
where $\mathcal{L}_{f}$ and $\mathcal{L}_{b}$ correspond to the losses for head and gesture motion, respectively. Further details of the reconstruction, velocity, and acceleration terms can be expanded based on the specific metrics used for evaluation. The total loss is computed as
@@
where $\lambda_{f}=3$ and $\lambda_{b}=1$ are weighting factors for the head and gesture motion losses. These weights control the relative contribution of each motion type to the overall optimization process.
@@ III-C Motion-Driven Detail Refinement Video Generate
Conditional Input Preparation. The video generation process is initiated by preparing the conditional inputs. As depicted in Fig. 5, the training inputs consist of the audio signal $a$, the original video $V$, the pose video $V_{P}$, and three types of motion masks: $V^{f+h}_{M}$, $V_{M}^{l}$, and $V_{M}^{b}$. These inputs collectively provide the necessary information to guide the generation of the final video output. We extract a 12-frame video clip from the training set to represent a single step of the model, with the first frame of the clip designated as the reference image $I_{0}$. This reference image serves as the foundational visual context for subsequent processing and generation within the model.
The CLIP [42] image encoder $E_{\text{clip}}$ extracts semantic features $Z_{text}$ from the text associated with the reference image $I_{0}$, while the autoencoder $E_{\text{vae}}$ [43, 44] encodes the implicit visual representation of the image. For the audio input $a$, we utilize $E_{a}$ [45, 46] to extract the corresponding audio features $f_{a}$. For the pose video $V_{P}$, a specially designed pose encoder [47] is employed to extract pose features, denoted as $Z_{\text{pose}}$. In addition, motion masks $V_{M}=\{V_{M}^{f+h},V_{M}^{l},V_{M}^{b}\}$ of different sizes are generated by Gaussian blurring and resizing operations. These masks are then fed into different layers of the model. All extracted features are temporally aligned to ensure that the generated video is synchronized with the input audio.
Figure 6: Qualitative Comparison of Upper Body Speech-Driven Video Generation Models on the PATS Dataset. This figure compares the performance of different speech-driven video generation models (MM-Diffusion[2], EDGE + Animate-anyone [48, 47], EDGE + hallo[48, 21] and Ours MMGT) in generating realistic upper-body movements, speech synchronization, and maintaining video quality.
Motion Masked Hierarchical Audio Attention. The Motion Mask-based Hierarchical Audio Attention (MM-HAA) module is a critical component in the second stage of our framework, designed to refine motion features by dynamically aligning spatial details, motion masks, and audio features. Unlike fixed-mask methods employed in prior works [21], MM-HAA leverages motion motion masks, $V_{M}^{f+h}$, $V_{M}^{l}$, and $V_{M}^{b}$ which are derived from keypoints generated in the first stage. These masks adaptively focus attention on specific regions, with $V_{M}^{f+h}$ focusing on face and gestures, $V_{M}^{l}$ focusing on lip movements, and $V_{M}^{b}$ focusing on the background area. Using a mechanism of cross-attention, MM-HAA aligns intermediate hidden states $Z_{CA}$ with specific temporal and frequency audio characteristics $f_{a}$, ensuring spatial-temporal synchronization. The key operation in MM-HAA is modeled as
@@
where $Q_{Z_{CA}}=W_{q}Z_{CA}$, $K_{f_{a}}=W_{k}f_{a}$, and $V_{f_{a}}=W_{v}f_{a}$ are the query, key, and value matrices, respectively, and $d_{k}$ represents the SMGA factor for stability. MM-HAA further integrates these aligned features with the motion masks to enhance region-specific refinement. The masked hidden states are computed as
$Z_{f+h}^{\prime}$ $=Z_{CA}\odot V_{M}^{f+h},$ (20)
$Z_{l}^{\prime}$ $=Z_{CA}\odot V_{M}^{l},$ (21)
@@
$Z_{\text{MM-HAA}}=\text{Adapter}_{f+h}(Z_{f+h}^{\prime})$ $+\text{Adapter}_{l}(Z_{l}^{\prime})$ (23)
@@
This combined output $Z_{\text{MM-HAA}}$ is passed to the Temporal Attention (TA) layers in the Denoising UNet, enabling further temporal refinement. The hierarchical design of MM-HAA ensures precise integration of motion masks, audio features, and spatial details across frames. By combining cross-attention and mask-based refinement, the module aligns audio-driven dynamics with spatially localized regions, effectively addressing the limitations of fixed-mask methods. Experimental results demonstrate that MM-HAA significantly improves realism, synchronization, and expressiveness in generated co-speech videos. Its contributions to temporal coherence and region-specific enhancement are essential to the overall success of the proposed framework.
Figure 7: Qualitative Comparison of Pose-driven Video Generation. This figure compares videos generated by our method and Animate-anyone [47] in a pose-driven task. In contrast, our method achieves high-quality generation with accurate and consistent hand movements.
Figure 8: Lips Synchronization with Audio. A comparison of lip synchronization with audio across Ground Truth, MM-Diffusion[2], EDGE + Animate-anyone [48, 47], EDGE + Hallo[48, 21], our MMGT model, and our ablation model (MMGT w/o SMGA).
@@
2:  $x_{0}\leftarrow E_{K}(V)=\{p^{0},...,p^{L}\}\triangleright\mathbb{R}^{L\times C}$
3:  $x_{r}\leftarrow repeat((E_{K}(I_{0})))=\{p^{0},...,p^{0}\}\triangleright\mathbb{R}^{L\times C}$
4:  $m_{body}=[0,1]\in\mathbb{R}^{L\times C_{b}}$
5:  $M=\{m_{body},1-m_{body}\}$
6:  $x_{t}\leftarrow concat(x_{0}+\mathcal{N}(0,\sigma^{2}),x_{r})$
7:  $\mathbf{c}_{1}=\{M,x_{t},E_{a}(a)\}$
@@
9:  $z_{0}\leftarrow E_{vae}(V)\triangleright\mathbb{R}^{L\times C\times H\times W}$
10:  $z_{t}\leftarrow z_{0}+\mathcal{N}(0,\sigma^{2})$
11:  $V_{M}=G_{M}(x_{0})$based on Algorithm.1}
12:  $V_{P}=Visualization(x_{0})$
13:  $\mathbf{c}_{2}=\{I,V_{M},V_{P},E_{a}(a)\}$
@@
21:     $=\theta_{2}-\nabla_{\theta_{2}}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta_{2}}\left(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\mathbf{c}_{2},t\right)\right\|^{2}$
@@
23:  $id_{\min}\leftarrow\text{Argmin}([\text{rmse}_{i}])$the $id$ of the min values}
24:  $\theta_{1}^{*}\leftarrow[\theta_{1}^{i}](id_{\min}),\theta_{2}^{*}\leftarrow[\theta_{2}^{i}](id_{\min})$

### v2

@@ III-B Pose Video and Motion Mask Generation
To introduce variations for motion modeling, Gaussian noise is added to the ground-truth pose $x_{0}=P_{\text{GT}}\in\mathbb{R}^{T\times C_{p}\times 3}$ to obtain noisy motion representations,
$x^{\prime}_{t}=x_{0}+\mathcal{N}(0,\sigma^{2}),$ (4)
where $x^{\prime}_{t}$ represents the noisy pose at time step $t$. The initial pose $p^{(0)}$ is repeated along the temporal dimension and added to $x^{\prime}_{t}$ obtain the input pose $x_{t}$,
$x_{t}=x^{\prime}_{t}+\mathcal{R}(p^{(0)}),$ (5)
where $\mathcal{R}(\cdot)$ denotes the replication of N copies of the delayed time dimension. The input pose $x_{t}$ is categorized into face spatial $C_{f}$ and body spatial $C_{b}$ as
@@
To apply masking to $x_{t}$ along the feature dimension $C$, we create the face spatial mask $M^{f}$ as
@@
Figure 6: Comparison of Algorithm 1 and Semantic Segmentation for Mask Generation. We further show a comparison of the stability of different mask generation methods when body parts overlap.
Next, the complementary body spatial mask $M^{b}$ is obtained as $M^{b}=1-M^{f}$. These spatial mask are then applied to the input $x_{t}$ to isolate the respective feature representations, as defined below
@@
where $x_{t}^{f}$ and $x_{t}^{b}$ correspond to the spatial features for head and body pose, respectively. Distinct encoders are employed to process these input conditions, transforming them into separate feature embeddings
@@
where $f_{a}$ represents audio feature embedding, with $E_{a}$ being a pre-trained audio encoder [17], while $E_{f}$ and $E_{b}$ are encoders for the face and body pose features, $f_{f}$ and $f_{b}$, respectively. Finally, the output of the first stage is expressed as
@@
where $t$ represents the temporal embedding. Subsequently, these features are fed into the Spatial Mask-Guided Audio2Pose (SMGA) Decoder, denoted as $D_{\text{SMGA}}$, which decodes them to produce a sequence of overlapping speech poses $\hat{x}_{0}$. The generated pose sequence $\hat{x}_{0}$ is then visualized as the pose video $V_{P}$.
Motion Mask Generation. As shown in Algorithm 1, we introduces a new way to create motion masks based on pose sequence, represented as $P\in\mathbb{R}^{T\times C_{p}\times 3}$. Here, $C_{p}=C_{f}+C_{h}+C_{l}$ corresponds to face, hand gestures, and lips, respectively. In each frame of a video sequence, our algorithm transforms normalized keypoint coordinates into pixel coordinates and calculates bounding boxes for specified regions. These bounding boxes are then utilized to construct binary masks, setting active regions to a value of 255. The generated masks include $V^{f}_{M}$ for face, $V_{M}^{l}$ for limb movements, and $V_{M}^{h}$ for hand gestures. Notably, as the head and hand masks are non-overlapping, they are combined into a single mask, denoted as $V_{M}^{f+h}$, and we set the background mask to $V_{M}^{b}=(255-V_{M}^{f+h})$. In the inference process of our framework, the generation process of the motion mask is shown in Fig. 6(a). Even in cases of limb occlusion or keypoints overlap, as shown in the green circle in Fig. 6(c)(d), the bounding boxes generated from different keypoints will only create overlapping areas instead of gaps, thus naturally ensuring full coverage of the motion region. It is superior to SAM2.1[39] in terms of stability of mask generation and is better suited to our framework.
@@
0: Pre-pose $\hat{P}\in\mathbb{R}^{N\times{C_{p}}},{C_{p}}=(C_{f}+C_{h}+C_{l})$
@@
SMGA Decoder Structure. The $D_{SMGA}$ is carefully designed to capture the temporal and semantic relationships between pose sequence and audio input. As shown in Fig. 5, its architecture consists of two motion blocks and a merging module. The two motion blocks process the input features $f_{f}$ or $f_{b}$ that represent different body parts separately through a combination of self-attention, FiLM[36, 47], and cross-attention mechanisms to capture the correlation between different body parts feature $f$ and audio feature $f_{a}$. Among them, the self-attention mechanism is used to capture the correlation of the pose sequence itself, and the calculation formula is as
Then, the cross-attention mechanism in different motion modules will capture the different correlations between different body parts $f$ and audio feature $f_{a}$. The specific formulas are as
where $Q_{f}$ is derived from pose sequence ($f_{f}$ or $f_{b}$), and $K_{f_{a}}$, $V_{f_{a}}$ are derived from audio features $f_{a}$. Through this distraction mechanism, we achieve coherent and coordinated speech actions for different body regions.
Next, inspired by EDGE [47], the joint representation is modulated by a FiLM block [36]. FiLM provides only affine modulation in the audio condition to enhance model representation. The diffusion time-step embedding $t$ with the audio feature $f_{a}$ after a linear layer MLP:
$x_{i}=f_{a}+\mathrm{MLP}(t),$ (13)
The FiLM block predicts the feature-wise scaling $\gamma_{i,c}$ and bias terms $\beta_{i,c}$:
$\gamma_{i,c}=f_{c}(x_{i}),\quad\beta_{i,c}=h_{c}(x_{i}),$ (14)
where $f_{c}(\cdot)$ and $h_{c}(\cdot)$ are learnable functions. FiLM then applies channel-wise affine transformations to three motion branches, $F_{i,c}$ with $c\in\{\text{face},\text{body},\text{merge}\}$:
$\operatorname{FiLM}(F_{i,c})=\gamma_{i,c}\,F_{i,c}+\beta_{i,c},$ (15)
where Face block injects $\gamma_{i,c},\beta_{i,c}$ to drive audio-synchronized face. Body block injects $\gamma_{i,c},\beta_{i,c}$ to synchronize gesture rhythms with the soundtrack. Merge block concatenates face and body features at the decoder’s end and applies a final FiLM gating.
Figure 7: The details of inputs during model training.
Loss Function. The training process optimizes the SMGA network using a multi-component loss function to ensure accurate pose generation. The reconstruction loss is defined as
@@
where $x_{t}$ and $\hat{x}_{t}$ represent the ground-truth and predicted pose sequence, respectively. The velocity loss is
$\mathcal{L}_{\text{vel}}=\frac{1}{T-1}\sum_{t=1}^{T-1}\|(x_{t+1}-x_{t})-(\hat{x}_{t+1}-\hat{x}_{t})\|_{2}^{2},$ (17)
@@
$\quad-\left(\hat{x}_{t+2}-2\hat{x}_{t+1}+\hat{x}_{t}\right)\Big\|_{2}^{2},$ (18)
Each loss term for head and body pose can be defined as
$\mathcal{L}_{f}=\mathcal{L}_{\text{rec}}^{f}+\mathcal{L}_{\text{vel}}^{f}+\mathcal{L}_{\text{acc}}^{f},\quad\mathcal{L}_{b}=\mathcal{L}_{\text{rec}}^{b}+\mathcal{L}_{\text{vel}}^{b}+\mathcal{L}_{\text{acc}}^{b},$ (19)
where $\mathcal{L}_{f}$ and $\mathcal{L}_{b}$ correspond to the losses for head and body pose, respectively. Further details of the reconstruction, velocity, and acceleration terms can be expanded based on the specific metrics used for evaluation. The total loss is computed as
@@
through sensitivity analysis, as shown in Table. IV, we selected $\lambda_{f}=3$ and $\lambda_{b}=1$ are weighting factors for the head and body pose losses. These weights control the relative contribution of each pose type to the overall optimization process.
@@ III-C Motion-Driven Detail Refinement Video Generate
Conditional Input Preparation. The video generation process begins with preparing the conditional inputs. As shown in Fig. 7, the training inputs consist of audio signal $a$, the original video $V$, the pose video $V_{P}$, and three types of motion masks: $V^{f+h}_{M}$, $V_{M}^{l}$, and $V_{M}^{b}$. These inputs collectively provide the necessary information to guide the generation of the final video output. We extract a 12-frame video clip from the training set to represent a single step of the model, with the first frame of the clip designated as the reference image $I_{0}$. This reference image provides the foundational visual context for subsequent processing and generation within the model.
The CLIP [38] image encoder $E_{\text{clip}}$ extracts semantic features $Z_{text}$ from the text associated with the reference image $I_{0}$, while the autoencoder $E_{\text{vae}}$ [23, 49] encodes the implicit visual representation of the image. For audio input $a$, we utilize $E_{a}$ [5, 10] to extract the corresponding audio features $f_{a}$. For the pose video $V_{P}$, a specially designed pose encoder [20] is employed to extract pose features, denoted as $Z_{\text{pose}}$. Additionally, motion masks $V_{M}=\{V_{M}^{f+h},V_{M}^{l},V_{M}^{b}\}$ of different sizes are generated using Gaussian blurring and resizing operations. These masks are then fed into different layers of the model. All extracted features are temporally aligned to ensure the generated video is synchronized with the input audio.
Figure 8: Qualitative Comparison of Upper Body Co-speech Video Generation Models on the PATS Dataset. This figure compares the performance of different co-speech video generation models ($M$-Diffusion[43], EDGE + Animate-anyone (EDGE + Animate) [47, 20], EDGE + Hallo[47, 54], DiffGest + MimicMo[63, 60] and Ours MMGT) in generating realistic upper-body movements, speech synchronization, and maintaining video quality
Motion Masked Hierarchical Audio Attention. The Motion Mask-based Hierarchical Audio Attention (MM-HAA) module is a vital part of the second stage of our framework, designed to refine motion features by dynamically aligning spatial details, motion masks, and audio features. Unlike fixed-mask methods used in previous works [54], MM-HAA uses motion masks, $V_{M}^{f+h}$, $V_{M}^{l}$, and $V_{M}^{b}$, which are derived from the pose sequence generated in the first stage. These masks adaptively focus attention on specific regions, with $V_{M}^{f+h}$ focusing on face and hand gestures, $V_{M}^{l}$ focusing on lip movements, and $V_{M}^{b}$ focusing on the background area. Through a cross-attention mechanism, MM-HAA aligns intermediate hidden states $Z_{CA}$ with specific temporal and frequency audio features $f_{a}$, ensuring spatial-temporal synchronization. The main operation in MM-HAA is defined as
@@
where $Q_{Z_{CA}}=W_{q}Z_{CA}$, $K_{f_{a}}=W_{k}f_{a}$, and $V_{f_{a}}=W_{v}f_{a}$ represent the query, key, and value matrices, respectively, and $d_{k}$ indicates the stability factor. MM-HAA also combines these aligned features with motion masks to improve region-specific refinement. The masked hidden states are computed as
$Z_{f+h}^{\prime}$ $=Z_{CA}\odot V_{M}^{f+h},$
$Z_{l}^{\prime}$ $=Z_{CA}\odot V_{M}^{l},$
@@
$Z_{\text{MM-HAA}}=\text{Adapter}_{f+h}(Z_{f+h}^{\prime})$ $+\text{Adapter}_{l}(Z_{l}^{\prime})$
@@
This combined output $Z_{\text{MM-HAA}}$ is fed into the Temporal Attention (TA) layers in the Denoising UNet, enabling further temporal refinement. The hierarchical design of MM-HAA ensures accurate integration of motion masks, audio features, and spatial details across frames. By combining cross-attention with mask-based refinement, the module aligns audio-driven dynamics with spatially localized regions, effectively addressing the limitations of fixed-mask methods. Experimental results show that MM-HAA significantly enhances realism, synchronization, and expressiveness in generated co-speech videos. Its contributions to temporal coherence and region-specific enhancement are crucial to the overall success of the proposed framework.
Figure 9: Qualitative Comparison of Pose-driven Video Generation. This figure compares videos generated by Animate-anyone [20], MimicMo [60]and MMGT(Ours) under the same pose-driven task.
TABLE I: Comparison of computational overhead of two training processes in the second stage of different methods.
Model (Processes I/Processes II) Params (GB) Train GPU-hours (h) Peak VRAM (GB) Inference(s/per-Clip) Batch size-per GPU Step
Animate-anyone[20] 6.3/1.69 7/13 41.4/35.88 151 2/1 29850/32500
MimicMo[60] 5.68 28 47.34 133 1 32500
Hallo[54] 6.3/2.80 7/28 41.4/16.72 240 2/1 29850/32500
MMGT(Ours) 6.3/2.17 7/15 41.4/41.27 224 2/1 29850/32500
Figure 10: Lips Synchronization with Audio. A comparison of lip synchronization with audio across Ground Truth, $M$-Diffusion[43], EDGE + Animate-anyone [47, 20], EDGE + Hallo[47, 54], DiffGest + MimicMo[63, 60] and our MMGT model.
@@
[…]
@@
[…]
@@
21:  $\theta_{2}-\nabla_{\theta_{2}}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta_{2}}\left(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\mathbf{c}_{2},t\right)\right\|^{2}$
@@
[…]